# Variational Bayesian Bow tie Neural Networks with Shrinkage

### v2

@@ C.1 Initialization schemes
Initialization plays an important role in the ability of Bayesian inference algorithms to effectively approximate the posterior. This is especially true in variational schemes for complex posteriors (such as for BNNs), which are only guaranteed to converge to local optimum. We design two possible variations of random yet effective initialization schemes. To simplify the exposition, we describe the procedure in the case of Inverse Gamma shrinkage priors, for which $\lambda=0$ and the selection of the scale parameters $\delta$ determines the level of shrinkage. Note that during the training step, we employ the expectation-maximization algorithm to set an optimal $\delta_{\text{glob}}$, whilst the value of $\delta_{loc,l}$ remains the fixed. To encourage more shrinkage for larger depth, we assume $\delta_{\text{glob}}\propto 1/\sqrt{L}$, and to encourage shrinkage for larger width set $\delta_{\text{loc},l}\propto 1/\sqrt{D_{l}}$. Given specified values of $\nu_{\text{loc}},\nu_{\text{glob}},\delta_{\text{loc}},\delta_{\text{glob}},\alpha^{h}_{0},\alpha_{0},\beta^{h}_{0},\beta_{0}$, we first re-scale the shrinkage parameters to scale appropriately
@@
  1. 1.
  2. 2.
  3. 3.
  4. 4.
    Variational parameters for $\bm{\tau},\bm{\psi}$:
  5. 5.
  6. 6.
C.2 Implementation details
@@
where $\mathbf{z}_{n,0}=\mathbf{x}_{n},\;\gamma\sim\text{IG}(2,1)\text{ and }l=1,\ldots,L,\;d=1,\ldots D_{l},\;d^{\prime}=1,\ldots,D_{l-1}.$ The choice of $\sigma_{y},\sigma_{W}$ and $\sigma_{b}$ is made in accordance with $\alpha_{0},s_{0}$ and $\delta_{\text{loc},l}$, respectively. For experiments with SVI we use Adam optimizer with learning rate set to 0.001 and maximum number of iterations varying from 5000 to 20000 depending on the dataset and depth of the network. Additionally, we consider the Bayes by Backprop model of Blundell et al. (2015) and adapt its Pytorch implementation from the publicly available repository Javier (2019). For all experiments with BBB we set the learning rate to 0.01 and maximum number of epochs varies from 500 to 1000.
Suppose that the data on which we evaluate the predictive performance consists of $N$ points and the true target is $\mathbf{y}^{*}$, then recorded evaluation metrics are RMSE, NLL and EC and are computed as follows:
@@
where the predicted observations are $\mathbf{y}^{o}$ and the corresponding quantiles are denoted as $q^{o}$. When computing quantiles to obtain empirical coverage and illustrating the uncertainty in Section 4 and below in Section C.3, we rely on the Gaussian approximation.
C.3 Supplementary material to the diabetes example.
Figure 10 supplements Table 2 and the diabetes example in Section 4.2. Here, in the case of VBNN, BBB and SVI models we provide the uncertainty of the observations and in the case of the LassoCV we provide residual standard deviation. Additionally, we illustrate the sparse prediction and the uncertainty obtained from sparse weights of the VBNN, which largely coincide with the original prediction and uncertainty estimates. Whilst the coverage estimates for observations of VBNN and BBB are comparable, the SVI underestimates the uncertainty and provides a dramatically lower coverage for observations.
@@
C.4 Supplementary information on the datasets
@@
Boston housing Harrison and Rubinfeld (1978): $n=506,p=13$, the predictors are per capita crime rate by town, the proportion of residential land zoned for lots over 25,000 sq.ft., the proportion of non-retail business acres per town, Charles River dummy variable, nitrite oxides concentration, average number of rooms per dwelling, the proportion of owner-occupied, units built before 1940, weighted distances to five Boston employment centres, index of accessibility to radial highways, full-value property-tax rate, the pupil-teacher ratio by town, the quantitative measure of systemic racism as a factor in house pricing, lower status of the population; the response of interest is the median value of owner-occupied homes. The Boston housing dataset is among the most popular pip available datasets, and with respect to variable selection it was considered in e.g. Schäfer and Chopin (2013).
@@
Energy Tsanas and Xifara (2012): $n=768,p=8$, the predictors are relative compactness, surface area, wall area, roof area, overall height, orientation, glazing area, and glazing area distribution, and the task is to predict the heating load of residential buildings.
@@
Yacht dynamics J. et al. (2013): $n=308,p=6$, the predictors are long position, prismatic coefficient, length-displacement ratio, bean-draught ratio length-bean ratio and froude number, and the task is to model the residuary resistance per unit weight of displacement for a yacht hull.
@@
Concrete compressive strength Yeh (2007): $n=1030,p=8$, the predictors are cement, furnace slag, fly ash, water, superplasticizer, coarse aggregate, fine aggregate and the age of testing, and the response variable is the compressive strength of concrete. This is also considered from the variable selection perspective in several works including Schäfer and Chopin (2013); Griffin (2024).
@@
Concrete slump test Yeh (2009): $n=103,p=7$, the predictors are concrete ingredients, namely cement, furnace slag, fly ash, water, superplasticizer, coarse aggregate, and fine aggregate and the task is to predict slump of concrete.
Appendix D Review of relevant distributions
D.1 Generalized Inverse Gaussian
@@
    where $\nu<0$ and $\delta>0$. This can also be re-written in terms of the more standard parametrization of the IG:
@@
    Note that if $w\sim\text{N}(0,\tau)$ and $\tau\sim\text{Gam}(1,\beta)$, this implies a marginal Laplace prior on $w$ (i.e. Bayesian Lasso Park and Casella (2008)) with scale $s=1/\sqrt{2\beta}=1/\lambda$.
@@
    Note that if $w\sim\text{N}(0,\tau)$ and $\tau\sim\text{IGaus}(\alpha,\beta)$, the marginal distribution is of the form Caron and Doucet (2008):
@@
D.2 EM update for different cases of global-local priors

### v3

@@ D.1 Initialization schemes
Initialization plays an important role in the ability of Bayesian inference algorithms to effectively approximate the posterior. This is especially true in variational schemes for complex posteriors (such as for BNNs), which are only guaranteed to converge to local optimum. We design two possible variations of random yet effective initialization schemes. To simplify the exposition, we describe the procedure in the case of Inverse Gamma shrinkage priors, for which $\lambda=0$ and the selection of the scale parameters $\delta$ determines the level of shrinkage. Note that during the training step, we employ the expectation-maximization algorithm to set an optimal $\delta_{\text{glob}}$, whilst the value of $\delta_{loc,l}$ remains fixed. To encourage more shrinkage for larger depth, we assume $\delta_{\text{glob}}\propto 1/\sqrt{L}$, and to encourage shrinkage for larger width set $\delta_{\text{loc},l}\propto 1/\sqrt{D_{l}}$. Given specified values of $\nu_{\text{loc}},\nu_{\text{glob}},\delta_{\text{loc}},\delta_{\text{glob}},\alpha^{h}_{0},\alpha_{0},\beta^{h}_{0},\beta_{0}$, we first re-scale the shrinkage parameters to scale appropriately
@@
D.2 Implementation details
@@
where $\mathbf{z}_{n,0}=\mathbf{x}_{n},\;\gamma\sim\text{IG}(2,1)\text{ and }l=1,\ldots,L,\;d=1,\ldots D_{l},\;d^{\prime}=1,\ldots,D_{l-1}.$ The choice of $\sigma_{y},\sigma_{W}$ and $\sigma_{b}$ is made in accordance with $\alpha_{0},s_{0}$ and $\delta_{\text{loc},l}$, respectively. For experiments with mfVI we use Adam optimizer with learning rate set to 0.001 and maximum number of iterations varying from 5000 to 20000 depending on the dataset and depth of the network. Additionally, we consider the Bayes by Backprop model of (Blundell et al., 2015) and adapt its Pytorch implementation from the publicly available repository (Javier, 2019). For all experiments with BBB we set the learning rate to 0.01 and maximum number of epochs varies from 500 to 1000.
In all examples, we normalize the input but do not re-scale the output. Suppose that the data on which we evaluate the predictive performance consists of $N$ points and the true target is $\mathbf{y}^{*}$, then recorded evaluation metrics are RMSE, NLL and EC and are computed as follows:
@@
where the predicted observations are $\mathbf{y}^{o}$ and the corresponding quantiles are denoted as $q^{o}$. When computing quantiles to obtain empirical coverage and illustrating the uncertainty in Section 4 and below in Section D.3, we rely on the Gaussian approximation.
D.3 Supplementary material to the diabetes example.
Figure 12 supplements Table 3 and the diabetes example in Section 4.2. Here, in the case of VBNN, BBB and mfVI models we provide the uncertainty of the observations and in the case of the LassoCV we provide residual standard deviation. Additionally, we illustrate the sparse prediction and the uncertainty obtained from sparse weights of the VBNN, which largely coincide with the original prediction and uncertainty estimates. Whilst the coverage estimates for observations of VBNN and BBB are comparable, the mfVI underestimates the uncertainty and provides a dramatically lower coverage for observations.
@@
D.4 Supplementary information on the datasets
@@
Boston housing (Harrison and Rubinfeld, 1978): $n=506,p=13$, the predictors are per capita crime rate by town, the proportion of residential land zoned for lots over 25,000 sq.ft., the proportion of non-retail business acres per town, Charles River dummy variable, nitrite oxides concentration, average number of rooms per dwelling, the proportion of owner-occupied, units built before 1940, weighted distances to five Boston employment centres, index of accessibility to radial highways, full-value property-tax rate, the pupil-teacher ratio by town, the quantitative measure of systemic racism as a factor in house pricing, lower status of the population; the response of interest is the median value of owner-occupied homes. The Boston housing dataset is among the most popular pip available datasets, and with respect to variable selection, it was considered in e.g. (Schäfer and Chopin, 2013).
@@
Energy (Tsanas and Xifara, 2012): $n=768,p=8$, the predictors are relative compactness, surface area, wall area, roof area, overall height, orientation, glazing area, and glazing area distribution, and the task is to predict the heating load of residential buildings.
@@
Yacht dynamics (J. et al., 2013): $n=308,p=6$, the predictors are long position, prismatic coefficient, length-displacement ratio, bean-draught ratio, length-bean ratio and froude number, and the task is to model the residuary resistance per unit weight of displacement for a yacht hull.
@@
Concrete compressive strength (Yeh, 2007): $n=1030,p=8$, the predictors are cement, furnace slag, fly ash, water, superplasticizer, coarse aggregate, fine aggregate and the age of testing, and the response variable is the compressive strength of concrete. This is also considered from the variable selection perspective in several works, including (Schäfer and Chopin, 2013; Griffin, 2024).
@@
Concrete slump test (Yeh, 2009): $n=103,p=7$, the predictors are concrete ingredients, namely cement, furnace slag, fly ash, water, superplasticizer, coarse aggregate, and fine aggregate, and the task is to predict the slump of concrete.
D.5 Supplementary material to the UCI datasets experiments
Table 4 supplements Figure 10, Figure 11 and the experiments described in Figure 10.
Table 4: RMSE, NLL and Coverage for UCI datasets.
Dataset
Metric Method Slump Yacht Boston Energy Concrete
RMSE 4SVIBNN $7.15\pm 1.5$ $3.57\pm.8$ $3.38\pm.9$ $1.62\pm.2$ $7.15\pm.6$
4VBNN $7.01\pm 1.2$ $1.23\pm.3$ $3.21\pm.6$ $1$2 $6.71\pm.6$
SVBNN $7.36\pm 1.62$ $5.57\pm 1.14$ $3.76\pm.92$ $2.41\pm.49$ $8.43\pm.75$
VBNN $7.37\pm 1.4$ $2.47\pm 1.1$ $3.47\pm.8$ $1.37\pm.3$ $7.67\pm.9$
GVBNN $7.64\pm 1.21$ $4.88\pm 2.66$ $4.02\pm.88$ $2.5\pm.42$ $7.84\pm.68$
mfVI $7.9\pm 1.7$ $61$36 $3.29\pm.6$ $2.27\pm.25$ 6.11±.5plus-or-minus6.11.56.$1$5
BBB $7.33\pm 1.94$ $1.45\pm.6$ $3.46\pm.96$ $2.65\pm.26$ $6.48\pm.61$
HMC $7.\pm 1.24$$56\pm.13$ $2.42\pm.47$$3\pm.07$ $4.01\pm.78$
HSBNN $6.41\pm 1.28$ $1.2\pm.23$ $2.92\pm.55$$6\pm.09$ $5.21\pm.56$
NLL 4SVBNN $3.42\pm 0.2$ $2.74\pm.2$ $2.61\pm.2$ $1.94\pm.1$ $3.38\pm.07$
4VBNN $3.39\pm.2$ $1.97\pm.2$ $2.57\pm.13$ $1.63\pm.16$ $3$06
SVBNN $3.47\pm.29$ $3.13\pm.27$ $2.78\pm.31$ $2.31\pm.22$ $3.56\pm 0.11$
VBNN $3.46\pm 0.2$ $2.25\pm 0.4$ $2.69\pm.26$ $1.75\pm.21$ $3.5\pm.12$
GVBNN $3.47\pm 0.17$ $2.86\pm.59$ $2.82\pm.22$ $2.35\pm.15$ $3.48\pm.09$
mfVI 3.77±.5plus-or-minus3.77.53.$7$5 $1.96\pm.09$ $2.61\pm.22$ $2.28\pm.09$ $3.57\pm.2$
BBB $6.23\pm 2.76$ $1.69\pm.14$ $2.47\pm.16$ $2.08\pm.15$ $3.21\pm 0.13$
HMC $3.41\pm.24$$87\pm.1$ $2.28\pm.18$$23\pm.44$ $2.74\pm.27$
HSBNN $5.02\pm 1.79$ $1.31\pm.18$ $5.01\pm 1.24$ $1.08\pm.2$ $4.32\pm.69$
Coverage 4SVBNN .$95\pm.06$$98\pm.03$$97\pm.03$$98\pm.01$$97\pm.01$
4VBNN .$94\pm.04$$9$02 .$97\pm.02$$9$0 .$97\pm.02$
SVBNN .$91\pm.08$$93\pm.02$$94\pm.03$$91\pm.04$$93\pm.03$
VBNN .$92\pm.06$$96\pm.01$$95\pm.03$$98\pm.02$$94\pm.02$
GVBNN .$96\pm.04$$95\pm.04$$96\pm.02$$93\pm.04$$95\pm.02$
mfVI .$78\pm.1$$96\pm.03$$96\pm.01$$95\pm.03$$8\pm.04$
BBB .$75\pm.12$ $1.\pm.0$$97\pm.02$$9$0 .$97\pm.02$
HMC .$9\pm.08$$98\pm.02$$96\pm.02$$95\pm.03$$94\pm.03$
HSBNN .$67\pm.14$$94\pm.05$ $0.62\pm.08$$89\pm.04$$71\pm.06$
D.6 Supplementary material to the SVI experiments
Figure 13: RMSE, NLL, EC of SVI compared to CAVI and plotted for various mini-batch sizes as a function of the forgetting rate $k$; the most right plot compares training times, where bar labels indicate the scale of computational gains.
[…]
Figure 14: The predictive means and uncertainty estimate as a function of the first coordinate for SVI, CAVI and ensembles of SVI approximations.
The computational gains obtained by utilising stochastic gradients and subsampling would motivate further research along the lines of model combination. Indeed, SVI makes ensembling techniques particularly appealing, in our experiment with a relatively small synthetic dataset, an ensemble of 5 SVI approximations with a mini-batch size of 10 performs comparable to CAVI while being more than 5 times faster than CAVI (see Figure 13 and Figure 14).
Appendix E Review of relevant distributions
E.1 Generalized Inverse Gaussian
@@
    where $\nu<0$ and $\delta>0$. This can also be rewritten in terms of the more standard parametrization of the IG:
@@
    Note that if $w\sim\text{N}(0,\tau)$ and $\tau\sim\text{Gam}(1,\beta)$, this implies a marginal Laplace prior on $w$ (i.e. Bayesian Lasso (Park and Casella, 2008)) with scale $s=1/\sqrt{2\beta}=1/\lambda$.
@@
    Note that if $w\sim\text{N}(0,\tau)$ and $\tau\sim\text{IGaus}(\alpha,\beta)$, the marginal distribution is of the form (Caron and Doucet, 2008):
@@
E.2 EM update for different cases of global-local priors